# How Fair is Your Diffusion Recommender Model?

### v1

@@ 4. Results and discussion
In this section, we aim to answer the following research question: “Is DiffRec a fair recommender system compared to the state-of-the-art?”. To this end, we conduct a twofold analysis. First, we investigate the accuracy and fairness performance separately, and then we measure the trade-off among performance metrics to examine whether DiffRec is suitable to provide both fair and useful recommendations.
@@
| Metrics | BPRMF | ItemkNN | NeuMF | LightGCN | UltraGCN | XSimGCL | EASE | MultiVAE | RecVAE | DiffRec | L-DiffRec | |
| ML1M | Recall () | 10.67 | 9.87 | 10.44 | 10.16 | 10.74 | 10.46 | 6.45 | 8.25 | 9.68 | 10.71 | 10.38 |
| nDCG () | 12.94 | 12.22 | 12.67 | 13.09 | 12.82 | 12.80 | 8.62 | 12.04 | 12.86 | 13.19 | 12.85 | |
| Recall () | 0.43 | 1.08 | 0.32 | 0.19 | 0.06 | 0.14 | 0.70 | 0.48 | 0.53 | 0.86 | 0.66 | |
| nDCG () | 1.57 | 1.98 | 1.54 | 1.43 | 1.37 | 1.34 | 1.66 | 2.43 | 2.07 | 2.03 | 2.25 | |
| APLT () | 12.47 | 8.01 | 11.75 | 8.11 | 13.73 | 10.50 | 9.00 | 2.63 | 5.38 | 6.18 | 10.81 | |
| Exp () | 86.29 | 91.18 | 87.36 | 91.32 | 84.98 | 88.78 | 90.54 | 97.02 | 94.10 | 93.67 | 87.26 | |
| FTKY | Recall () | 12.17 | 11.77 | 12.48 | 11.03 | 12.22 | 13.07 | 6.19 | 10.06 | 12.30 | 13.10 | 12.48 |
| nDCG () | 10.73 | 10.48 | 10.93 | 9.87 | 10.85 | 11.37 | 5.48 | 9.21 | 10.90 | 11.53 | 10.77 | |
| Recall () | 0.94 | 0.69 | 0.59 | 0.77 | 1.05 | 0.98 | 0.40 | 0.97 | 1.00 | 0.80 | 0.61 | |
| nDCG () | 0.22 | 0.19 | 0.09 | 0.47 | 0.29 | 0.23 | 0.10 | 0.59 | 0.29 | 0.22 | 0.02 | |
| APLT () | 9.93 | 18.27 | 10.20 | 7.42 | 5.36 | 11.06 | 6.32 | 3.72 | 7.27 | 10.75 | 16.97 | |
| Exp () | 89.27 | 78.91 | 89.39 | 91.87 | 94.32 | 88.25 | 93.56 | 95.75 | 92.03 | 88.57 | 80.85 |
4.1. Single-metric performance
@@
Table 3 shows the results of the selected baselines on ML1M and FTKY in terms of recommendation accuracy, i.e., Recall and nDCG, consumer fairness, i.e., Recall and nDCG, and provider fairness, i.e., APLT and Exp. We recall that all metrics are calculated on top-20 recommendation lists, and are expressed as percentages (%).
@@
On the recommendation accuracy level, the results are almost aligned with the findings from the related literature, where graph- and diffusion-based recommender systems generally dominate over the other baselines on both recommendation datasets. DiffRec shows slightly superior performance than its variant L-DiffRec, but the latter is more computationally lightweight. This gap in performance between the two DiffRec variants might be due to the diffusion process being conducted in the latent space by L-DiffRec. Specifically, the variational autoencoders employed by L-DiffRec compress the users’ preferences signals, which might lead to information loss and a non-perfect reconstruction in the decoding process. On the other hand, DiffRec forward and reverse processes fully capture the discrete preference information, without dealing with the potential loss of details caused by information bottleneck.
@@
Moving to the experiments on fairness, results exhibit distinct trends across the datasets. On ML1M, graph-based approaches such as UltraGCN and XSimGCL are able to provide high-utility and fair recommendations, while the diffusion models settle among the unfairest approaches across all the baselines. On FTKY, the fairest recommender systems are instead the generative ones, such as EASE and L-DiffRec, and the traditional ones, such as ItemKNN and NeuMF. The diverse trends in consumer fairness across ML1M and FTKY might derive from their different domains. Specifically, users’ interactions in the movie domain are driven by intrinsically individual preferences towards specific movie categories, such as those sharing the same actors or the same genre, whereas venue check-ins in FTKY might derive from goings out decided as groups, e.g., with friends or partners. Biases in movies’ preferences might be exacerbated by the RSs compared with venues’ preferences.
@@
The consumer unfairness exhibited by DiffRec on ML1M underscores that diffusion models are prone to increase the bias impact, as also emphasized by other studies in different domains. However, a common observation across the datasets is that the L-DiffRec variant reports fairer recommendations than DiffRec and the other family-related generative models, e.g., MultiVAE and RecVAE. Particularly, regarding APLT and Exp, L-DiffRec exhibits fairer outcomes on the provider side. This might be attributed to the intrinsic clustering operation of L-DiffRec, which shifts the learning process from following the common popularity bias to capturing more nuanced patterns across items’ categories.
4.2. Trade-off analysis
@@
As the three analyzed properties, i.e., accuracy, consumer, and provider fairness, represent different and possibly contrasting sides of the recommendation scenario, we extend our previous analysis by assessing the trade-off among them, as recently done by related studies (Anelli et al., 2023; Boratto et al., 2023b). Indeed, we believe the current analysis may suitably complement the findings from the single-metric study conducted above, motivating and explaining from other complementary perspectives the previous performance trends.
@@
The Kiviat diagrams in Figure 1 represent the performance for both datasets of five representative baselines, i.e., ItemkNN, XSimGCL, RecVAE, DiffRec, and L-DiffRec, on three metrics, i.e., nDCG, nDCG, and APLT, accounting for accuracy, consumer, and provider fairness. To provide an easier interpretation of the plots, we follow (Anelli et al., 2023) and normalize the metrics into [0, 1] through min-max normalization and rescale nDCG to adhere to the principle that “the higher the values the better the performance”.
@@
On the ML1M dataset, most of the findings from the single-metric analysis are generally confirmed. Indeed, we observe that XSimGCL is the most balanced baseline across all models, being capable of providing an optimal trade-off among all metrics. Conversely, ItemkNN, RecVAE, and DiffRec cannot generally increase one metric side without harming the other ones. Interestingly, though, while ItemkNN tends to maximize fairness over the accuracy, RecVAE and DiffRec, i.e., generative and diffusion-based models, do the opposite. This becomes especially evident on DiffRec, where the model is largely improving accuracy at the detriment of provider/consumer fairness. The exception to the trend is L-DiffRec, reaching the second-best trade-off soon after XSimGCL. In fact, on FTKY, L-DiffRec exhibits even the best trade-off compared to the selected baselines.
Thus, the findings from the trade-off analysis, with a more clarifying visualization, confirm the unfairness warnings raised in the single-metric study but, at the same time, suggest there might be room for improvement in diffusion-based recommendation toward more fairness-aware recommendations.
@@ 5. Conclusion and future work
This work settles as a preliminary attempt to rigorously assess the performance of diffusion-based recommender systems, specifically focusing on one of the pioneer approaches, DiffRec, and its variant L-DiffRec. The relevance of our study is emphasized by the machine learning literature that has recently observed that the intrinsic bias within the data may cause diffusion models to generate unfair outcomes. To this end, we decided to tailor our investigation to the performance of (L-)DiffRec under common fairness properties in recommender systems. Specifically, we conducted an extensive analysis involving two recommendation datasets (with sensitive information), nine state-of-the-art recommendation systems, and six recommendation metrics accounting for accuracy, consumer, and provider fairness. First, in a single-metrics evaluation, results generally outlined that DiffRec might exhibit unfairness issues; however, such issues are mitigated by the L-DiffRec variant, especially on the provider side. Then, a complementary trade-off analysis among accuracy and fairness properties confirmed that L-DiffRec might be the key to reaching optimal metrics balancing. We believe the take-home message is that diffusion models cannot be plugged into recommender systems as they are, since they would bring (and even exacerbate) the unfairness issues already observed in machine learning. However, with careful and tailored modifications to the rationale and models’ architecture, e.g., L-DiffRec, the issues might be properly addressed. To confirm our final assumption, future works will aim to extend our study to the other DiffRec variant, T-DiffRec, as well as other diffusion-based recommender systems in the literature.
@@ Acknowledgements.
Daniele Malitesta and Fragkiskos D. Malliaros acknowledge the support by ANR (French National Research Agency) under the JCJC project GraphIA (ANR-20-CE23-0009-01). The authors acknowledge Inria for the availability of high-performance computing resources.References
@@
- Anelli et al. (2022) Vito Walter Anelli, Alejandro Bellogín, Tommaso Di Noia, Dietmar Jannach, and Claudio Pomo. 2022. Top-N Recommendation Algorithms: A Quest for the State-of-the-Art. In UMAP. ACM, 121–131.
@@
- Anelli et al. (2023) Vito Walter Anelli, Yashar Deldjoo, Tommaso Di Noia, Daniele Malitesta, Vincenzo Paparella, and Claudio Pomo. 2023. Auditing Consumer- and Producer-Fairness in Graph Collaborative Filtering. In ECIR (1) (Lecture Notes in Computer Science, Vol. 13980). Springer, 33–48.
- Austin et al. (2021) Jacob Austin, Daniel D. Johnson, Jonathan Ho, Daniel Tarlow, and Rianne van den Berg. 2021. Structured Denoising Diffusion Models in Discrete State-Spaces. In NeurIPS. 17981–17993.
@@
- Bianchi et al. (2023) Federico Bianchi, Pratyusha Kalluri, Esin Durmus, Faisal Ladhak, Myra Cheng, Debora Nozza, Tatsunori Hashimoto, Dan Jurafsky, James Zou, and Aylin Caliskan. 2023. Easily Accessible Text-to-Image Generation Amplifies Demographic Stereotypes at Large Scale. In FAccT. ACM, 1493–1504.
@@
- Boratto et al. (2023a) Ludovico Boratto, Francesco Fabbri, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2023a. Counterfactual Graph Augmentation for Consumer Unfairness Mitigation in Recommender Systems. In CIKM. ACM, 3753–3757.
@@
- Boratto et al. (2024) Ludovico Boratto, Francesco Fabbri, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2024. Robustness in Fairness Against Edge-Level Perturbations in GNN-Based Recommendation. In ECIR (3) (Lecture Notes in Computer Science, Vol. 14610). Springer, 38–55.
@@
- Boratto et al. (2022) Ludovico Boratto, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2022. Consumer Fairness in Recommender Systems: Contextualizing Definitions and Mitigations. In ECIR (1) (Lecture Notes in Computer Science, Vol. 13185). Springer, 552–566.
@@
- Boratto et al. (2023b) Ludovico Boratto, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2023b. Practical perspectives of consumer fairness in recommendation. Inf. Process. Manag. 60, 2 (2023), 103208.
- Cao et al. (2024) Hanqun Cao, Cheng Tan, Zhangyang Gao, Yilun Xu, Guangyong Chen, Pheng-Ann Heng, and Stan Z. Li. 2024. A Survey on Generative Diffusion Models. IEEE Transactions on Knowledge and Data Engineering (2024), 1–20.
@@
- Cho et al. (2023) Jaemin Cho, Abhay Zala, and Mohit Bansal. 2023. DALL-EVAL: Probing the Reasoning Skills and Social Biases of Text-to-Image Generation Models. In ICCV. IEEE, 3020–3031.
@@
- Choi et al. (2024) Yujin Choi, Jinseong Park, Hoki Kim, Jaewook Lee, and Saerom Park. 2024. Fair Sampling in Diffusion Models through Switching Mechanism. In AAAI. AAAI Press, 21995–22003.
@@
- Friedrich et al. (2023) Felix Friedrich, Patrick Schramowski, Manuel Brack, Lukas Struppek, Dominik Hintersdorf, Sasha Luccioni, and Kristian Kersting. 2023. Fair Diffusion: Instructing Text-to-Image Generation Models on Fairness. CoRR abs/2302.10893 (2023).
@@
- Goodfellow et al. (2020) Ian J. Goodfellow, Jean Pouget-Abadie, Mehdi Mirza, Bing Xu, David Warde-Farley, Sherjil Ozair, Aaron C. Courville, and Yoshua Bengio. 2020. Generative adversarial networks. Commun. ACM 63, 11 (2020), 139–144.
@@
- Harper and Konstan (2016) F. Maxwell Harper and Joseph A. Konstan. 2016. The MovieLens Datasets: History and Context. ACM Trans. Interact. Intell. Syst. 5, 4 (2016), 19:1–19:19.
@@
- He et al. (2020) Xiangnan He, Kuan Deng, Xiang Wang, Yan Li, Yong-Dong Zhang, and Meng Wang. 2020. LightGCN: Simplifying and Powering Graph Convolution Network for Recommendation. In SIGIR. ACM, 639–648.
@@
- He et al. (2017) Xiangnan He, Lizi Liao, Hanwang Zhang, Liqiang Nie, Xia Hu, and Tat-Seng Chua. 2017. Neural Collaborative Filtering. In WWW. ACM, 173–182.
- Ho et al. (2020) Jonathan Ho, Ajay Jain, and Pieter Abbeel. 2020. Denoising Diffusion Probabilistic Models. In NeurIPS.

### v2

@@ 4. Results and Discussion
In this section, we aim to answer the following research question: “Is DiffRec a fair RS compared to state-of-the-art methods?”. To this end, we first investigate the recommendation utility and fairness dimensions separately, and then we inspect the trade-off among metrics from both dimensions to assess whether DiffRec can deliver recommendations that are both fair and effective.
@@
| Metrics | Random | Pop | BPRMF | ItemkNN | NeuMF | LightGCN | UltraGCN | XSimGCL | EASE | MultiVAE | RecVAE | DiffRec | L-DiffRec |
| Movielens-1M (ML1M) | |||||||||||||
| Recall () | 0.61 | 6.88 | 10.67 | 9.87 | 10.44 | 10.16 | 10.74 | 10.46 | 6.45 | 8.25 | 9.68 | 10.71 | 10.38 |
| nDCG () | 1.14 | 10.42 | 12.94 | 12.22 | 12.67 | 13.09 | 12.82 | 12.80 | 8.62 | 12.04 | 12.86 | 13.19 | 12.85 |
| Recall () | 0.01 | 0.82 | 0.43 | 1.08 | 0.32 | 0.19 | 0.06 | 0.14 | 0.70 | 0.48 | 0.53 | 0.86 | 0.66 |
| nDCG () | 0.19 | 2.54 | 1.57 | 1.98 | 1.54 | 1.43 | 1.37 | 1.34 | 1.66 | 2.43 | 2.07 | 2.03 | 2.25 |
| APLT () | 79.80 | 0.00 | 12.47 | 8.01 | 11.75 | 8.11 | 13.73 | 10.50 | 9.00 | 2.63 | 5.38 | 6.18 | 10.81 |
| Exp () | 0.12 | 100.00 | 86.29 | 91.18 | 87.36 | 91.32 | 84.98 | 88.78 | 90.54 | 97.02 | 94.10 | 93.67 | 87.26 |
| Fousquare Tokyo (FTKY) | |||||||||||||
| Recall () | 0.30 | 7.89 | 12.17 | 11.77 | 12.48 | 11.03 | 12.22 | 13.07 | 6.19 | 10.06 | 12.30 | 13.10 | 12.48 |
| nDCG () | 0.22 | 5.57 | 10.73 | 10.48 | 10.93 | 9.87 | 10.85 | 11.37 | 5.48 | 9.21 | 10.90 | 11.53 | 10.77 |
| Recall () | 0.03 | 0.95 | 0.94 | 0.69 | 0.59 | 0.77 | 1.05 | 0.98 | 0.40 | 0.97 | 1.00 | 0.80 | 0.61 |
| nDCG () | 0.01 | 0.85 | 0.22 | 0.19 | 0.09 | 0.47 | 0.29 | 0.23 | 0.10 | 0.59 | 0.29 | 0.22 | 0.02 |
| APLT () | 80.00 | 0.00 | 9.93 | 18.27 | 10.20 | 7.42 | 5.36 | 11.06 | 6.32 | 3.72 | 7.27 | 10.75 | 16.97 |
| Exp () | 0.09 | 100.00 | 89.27 | 78.91 | 89.39 | 91.87 | 94.32 | 88.25 | 93.56 | 95.75 | 92.03 | 88.57 | 80.85 |
4.1. Uni-Dimensional Analysis
@@
Table 3 shows the results of the selected baselines (along with reference models such as Random and Pop) on ML1M and FTKY in terms of recommendation utility (i.e., Recall and nDCG), consumer fairness (i.e., Recall and nDCG), and provider fairness (i.e., APLT and Exp). Note that each metric is computed on the top-20 recommendation lists and is expressed as a percentage (%).
@@
On the recommendation utility dimension, the results are closely aligned with those observed in the related literature, where graph- and diffusion-based RSs generally dominate over the other baselines on both recommendation datasets. DiffRec shows slightly superior performance than its variant L-DiffRec. However, it should be noted that the latter is more computationally lightweight. This gap between the two DiffRec variants might be due to the diffusion process being conducted in the latent space by L-DiffRec. The VAEs employed by L-DiffRec compress the users’ preferences signals, potentially leading to information loss and non-perfect reconstruction during decoding. Conversely, the DiffRec forward and reverse processes fully capture the discrete preference information, without dealing with the potential loss caused by information bottlenecks.
@@
Moving to the experiments on the fairness dimension, our results exhibit distinct trends across the datasets. On ML1M, graph-based approaches such as UltraGCN and XSimGCL can provide both highly effective and fair recommendations, while the diffusion-based models settle among the least fair approaches among the considered methods. On FTKY, the fairest RSs are instead the generative ones, such as EASE and L-DiffRec, and the traditional ones, such as ItemKNN and NeuMF. The diverse trends in consumer fairness across ML1M and FTKY might derive from their different domains. Specifically, users’ interactions in the movie domain are driven by intrinsically individual preferences towards specific movie categories, such as those sharing the same actors or the same genre. In contrast, venue check-ins in FTKY might derive from going out decided as groups, e.g., with friends or partners. Biases in movie preferences may be more strongly amplified by recommender systems compared to biases in venue preferences.
@@
The consumer unfairness exhibited by DiffRec on ML1M shows that diffusion models are prone to amplify biases, as also reported by other studies. However, a common observation across the datasets is that the L-DiffRec variant leads to fairer recommendations than DiffRec and the other considered generative models, e.g., MultiVAE and RecVAE. Particularly, regarding APLT and Exp, L-DiffRec exhibits fairer outcomes on the provider side. This might be attributed to the intrinsic clustering operation of L-DiffRec, which shifts the learning process from following the common popularity bias to capturing more nuanced patterns across items’ categories.
4.2. Multi-Dimensional Analysis
@@
As the three analyzed properties, i.e., recommendation utility, consumer fairness, and provider fairness, represent different and possibly contrasting dimensions of the recommendation scenario, we extend our previous analysis by assessing the trade-off among them, as done in related studies (Anelli et al., 2023; Boratto et al., 2023b). Indeed, we believe that this additional assessment may complement the findings from the uni-dimensional analysis, motivating and explaining from other complementary perspectives the previous performance trends.
@@
The Kiviat diagrams in Figure 1 represent the performance for both datasets of five representative baselines, i.e., ItemkNN, XSimGCL, RecVAE, DiffRec, and L-DiffRec, on three metrics, i.e., nDCG, nDCG, and APLT, accounting for utility, consumer, and provider fairness. To provide an easier interpretation of the plots, we follow (Anelli et al., 2023) and normalize the metrics into [0, 1] through min-max normalization and rescale nDCG to adhere to the principle that “the higher the values the better the performance”.
@@
On the ML1M dataset, the findings from the uni-dimensional analysis are generally confirmed. Indeed, we observe that XSimGCL is the most balanced baseline across all models, being capable of providing an optimal trade-off among all metrics. Conversely, ItemkNN, RecVAE, and DiffRec cannot generally increase one metric side without harming the other ones. Interestingly, though, while ItemkNN tends to maximize fairness over the utility, RecVAE and DiffRec, i.e., generative and diffusion-based models, do the opposite. This becomes especially evident on DiffRec, where the model is largely improving utility at the detriment of provider/consumer fairness. The exception to the trend is L-DiffRec, reaching the second-best trade-off soon after XSimGCL. In fact, on FTKY, L-DiffRec exhibits even the best trade-off compared to the selected baselines.
In summary, the findings from the multi-dimensional analysis confirm the unfairness warnings raised in the uni-dimensional analysis but, at the same time, suggest that diffusion-based methods may have room for improvement in achieving fairer recommendations.
@@ 5. Conclusion and Future Work
This work serves as a preliminary attempt to rigorously assess the performance of diffusion-based RSs, specifically focusing on one of the pioneer approaches, DiffRec, and its variant L-DiffRec. The relevance of our study is underscored by the ML literature, which has recently noted that the intrinsic bias within the data may lead to diffusion models generating unfair outcomes. To this end, we decided to tailor our investigation to the performance of (L-)DiffRec under common fairness properties in RSs. Specifically, we conducted an extensive analysis involving two recommendation datasets (with sensitive information), nine state-of-the-art RSs, and six recommendation metrics that account for accuracy, consumer, and provider fairness.
Results in two complementary evaluation settings, conducted on single metrics and in a trade-off scenario, empirically confirmed that the same bias amplification caused by diffusion models in ML seems to be advisable also in the recommendation domain (DiffRec). However, when adopting diffusion strategies carefully designed to tailor the recommendation scenario (L-DiffRec), the unfairness issue can be solved and even reverted. Thus, our empirical study suggests that diffusion models cannot be plugged into RSs as they are, since they would bring (and even exacerbate) the unfairness issues already observed in ML.
However, with careful and tailored modifications to the rationale and models’ architecture (e.g., operating in a clustered latent space, as done in L-DiffRec, to reduce noise sensitivity and group-level bias), the issues might be properly addressed.
To confirm our final assumption, future works will aim to extend our study to the other DiffRec variant, T-DiffRec, as well as other diffusion-based RSs (Walker et al., 2022; Hou et al., 2024; Choi et al., 2024). While trying to confirm the same observed (un)fairness trends as in this work, we will also devote future investigations to understanding the source of unfairness more carefully, for instance, by examining how the recommendation dataset is modified after the generation process of the diffusion model. Finally, on such theoretical and empirical validations, we plan to exploit the lessons learned from this empirical study and propose tailored strategies to mitigate (when present) the unfairness of diffusion-based recommendation systems.
@@ Acknowledgements.
Daniele Malitesta and Fragkiskos D. Malliaros acknowledge the support of the Innov4-ePiK project managed by the French National Research Agency under the 4th PIA, integrated into France2030 (ANR-23-RHUS-0002).References
@@
- Anelli et al. (2022) Vito Walter Anelli, Alejandro Bellogín, Tommaso Di Noia, Dietmar Jannach, and Claudio Pomo. 2022. Top-N Recommendation Algorithms: A Quest for the State-of-the-Art. In UMAP ’22: 30th ACM Conference on User Modeling, Adaptation and Personalization, Barcelona, Spain, July 4 - 7, 2022, Alejandro Bellogín, Ludovico Boratto, Olga C. Santos, Liliana Ardissono, and Bart P. Knijnenburg (Eds.). ACM, 121–131. https://doi.org/10.1145/3503252.3531292
@@
- Anelli et al. (2023) Vito Walter Anelli, Yashar Deldjoo, Tommaso Di Noia, Daniele Malitesta, Vincenzo Paparella, and Claudio Pomo. 2023. Auditing Consumer- and Producer-Fairness in Graph Collaborative Filtering. In Advances in Information Retrieval - 45th European Conference on Information Retrieval, ECIR 2023, Dublin, Ireland, April 2-6, 2023, Proceedings, Part I (Lecture Notes in Computer Science, Vol. 13980), Jaap Kamps, Lorraine Goeuriot, Fabio Crestani, Maria Maistro, Hideo Joho, Brian Davis, Cathal Gurrin, Udo Kruschwitz, and Annalina Caputo (Eds.). Springer, 33–48. https://doi.org/10.1007/978-3-031-28244-7_3
- Austin et al. (2021) Jacob Austin, Daniel D. Johnson, Jonathan Ho, Daniel Tarlow, and Rianne van den Berg. 2021. Structured Denoising Diffusion Models in Discrete State-Spaces. In Advances in Neural Information Processing Systems 34: Annual Conference on Neural Information Processing Systems 2021, NeurIPS 2021, December 6-14, 2021, virtual, Marc’Aurelio Ranzato, Alina Beygelzimer, Yann N. Dauphin, Percy Liang, and Jennifer Wortman Vaughan (Eds.). 17981–17993. https://proceedings.neurips.cc/paper/2021/hash/958c530554f78bcd8e97125b70e6973d-Abstract.html
- Balloccu et al. (2023) Giacomo Balloccu, Ludovico Boratto, Christian Cancedda, Gianni Fenu, and Mirko Marras. 2023. Knowledge is Power, Understanding is Impact: Utility and Beyond Goals, Explanation Quality, and Fairness in Path Reasoning Recommendation. In Advances in Information Retrieval - 45th European Conference on Information Retrieval, ECIR 2023, Dublin, Ireland, April 2-6, 2023, Proceedings, Part III (Lecture Notes in Computer Science, Vol. 13982), Jaap Kamps, Lorraine Goeuriot, Fabio Crestani, Maria Maistro, Hideo Joho, Brian Davis, Cathal Gurrin, Udo Kruschwitz, and Annalina Caputo (Eds.). Springer, 3–19. https://doi.org/10.1007/978-3-031-28241-6_1
@@
- Bianchi et al. (2023) Federico Bianchi, Pratyusha Kalluri, Esin Durmus, Faisal Ladhak, Myra Cheng, Debora Nozza, Tatsunori Hashimoto, Dan Jurafsky, James Zou, and Aylin Caliskan. 2023. Easily Accessible Text-to-Image Generation Amplifies Demographic Stereotypes at Large Scale. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency, FAccT 2023, Chicago, IL, USA, June 12-15, 2023. ACM, 1493–1504. https://doi.org/10.1145/3593013.3594095
@@
- Boratto et al. (2023a) Ludovico Boratto, Francesco Fabbri, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2023a. Counterfactual Graph Augmentation for Consumer Unfairness Mitigation in Recommender Systems. In Proceedings of the 32nd ACM International Conference on Information and Knowledge Management, CIKM 2023, Birmingham, United Kingdom, October 21-25, 2023, Ingo Frommholz, Frank Hopfgartner, Mark Lee, Michael Oakes, Mounia Lalmas, Min Zhang, and Rodrygo L. T. Santos (Eds.). ACM, 3753–3757. https://doi.org/10.1145/3583780.3615165
- Boratto et al. (2024a) Ludovico Boratto, Francesco Fabbri, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2024a. Fair Augmentation for Graph Collaborative Filtering. In Proceedings of the 18th ACM Conference on Recommender Systems, RecSys 2024, Bari, Italy, October 14-18, 2024, Tommaso Di Noia, Pasquale Lops, Thorsten Joachims, Katrien Verbert, Pablo Castells, Zhenhua Dong, and Ben London (Eds.). ACM, 158–168. https://doi.org/10.1145/3640457.3688064
@@
- Boratto et al. (2024b) Ludovico Boratto, Francesco Fabbri, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2024b. Robustness in Fairness Against Edge-Level Perturbations in GNN-Based Recommendation. In Advances in Information Retrieval - 46th European Conference on Information Retrieval, ECIR 2024, Glasgow, UK, March 24-28, 2024, Proceedings, Part III (Lecture Notes in Computer Science, Vol. 14610), Nazli Goharian, Nicola Tonellotto, Yulan He, Aldo Lipani, Graham McDonald, Craig Macdonald, and Iadh Ounis (Eds.). Springer, 38–55. https://doi.org/10.1007/978-3-031-56063-7_3
@@
- Boratto et al. (2022) Ludovico Boratto, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2022. Consumer Fairness in Recommender Systems: Contextualizing Definitions and Mitigations. In Advances in Information Retrieval - 44th European Conference on IR Research, ECIR 2022, Stavanger, Norway, April 10-14, 2022, Proceedings, Part I (Lecture Notes in Computer Science, Vol. 13185), Matthias Hagen, Suzan Verberne, Craig Macdonald, Christin Seifert, Krisztian Balog, Kjetil Nørvåg, and Vinay Setty (Eds.). Springer, 552–566. https://doi.org/10.1007/978-3-030-99736-6_37
@@
- Boratto et al. (2023b) Ludovico Boratto, Gianni Fenu, Mirko Marras, and Giacomo Medda. 2023b. Practical perspectives of consumer fairness in recommendation. Inf. Process. Manag. 60, 2 (2023), 103208. https://doi.org/10.1016/J.IPM.2022.103208
- Chen et al. (2023) Xiaohui Chen, Jiaxing He, Xu Han, and Liping Liu. 2023. Efficient and Degree-Guided Graph Generation via Discrete Diffusion Modeling. In International Conference on Machine Learning, ICML 2023, 23-29 July 2023, Honolulu, Hawaii, USA (Proceedings of Machine Learning Research, Vol. 202), Andreas Krause, Emma Brunskill, Kyunghyun Cho, Barbara Engelhardt, Sivan Sabato, and Jonathan Scarlett (Eds.). PMLR, 4585–4610. https://proceedings.mlr.press/v202/chen23k.html
@@
- Cho et al. (2023) Jaemin Cho, Abhay Zala, and Mohit Bansal. 2023. DALL-EVAL: Probing the Reasoning Skills and Social Biases of Text-to-Image Generation Models. In IEEE/CVF International Conference on Computer Vision, ICCV 2023, Paris, France, October 1-6, 2023. IEEE, 3020–3031. https://doi.org/10.1109/ICCV51070.2023.00283
@@
- Choi et al. (2024) Yujin Choi, Jinseong Park, Hoki Kim, Jaewook Lee, and Saerom Park. 2024. Fair Sampling in Diffusion Models through Switching Mechanism. In Thirty-Eighth AAAI Conference on Artificial Intelligence, AAAI 2024, Thirty-Sixth Conference on Innovative Applications of Artificial Intelligence, IAAI 2024, Fourteenth Symposium on Educational Advances in Artificial Intelligence, EAAI 2014, February 20-27, 2024, Vancouver, Canada, Michael J. Wooldridge, Jennifer G. Dy, and Sriraam Natarajan (Eds.). AAAI Press, 21995–22003. https://doi.org/10.1609/AAAI.V38I20.30202
- Dwork et al. (2012) Cynthia Dwork, Moritz Hardt, Toniann Pitassi, Omer Reingold, and Richard S. Zemel. 2012. Fairness through awareness. In Innovations in Theoretical Computer Science 2012, Cambridge, MA, USA, January 8-10, 2012, Shafi Goldwasser (Ed.). ACM, 214–226. https://doi.org/10.1145/2090236.2090255
- Feldman et al. (2015) Michael Feldman, Sorelle A. Friedler, John Moeller, Carlos Scheidegger, and Suresh Venkatasubramanian. 2015. Certifying and Removing Disparate Impact. In Proceedings of the 21th ACM SIGKDD International Conference on Knowledge Discovery and Data Mining, Sydney, NSW, Australia, August 10-13, 2015, Longbing Cao, Chengqi Zhang, Thorsten Joachims, Geoffrey I. Webb, Dragos D. Margineantu, and Graham Williams (Eds.). ACM, 259–268. https://doi.org/10.1145/2783258.2783311
@@
- Friedrich et al. (2023) Felix Friedrich, Patrick Schramowski, Manuel Brack, Lukas Struppek, Dominik Hintersdorf, Sasha Luccioni, and Kristian Kersting. 2023. Fair Diffusion: Instructing Text-to-Image Generation Models on Fairness. CoRR abs/2302.10893 (2023). https://doi.org/10.48550/ARXIV.2302.10893 arXiv:2302.10893
@@
- Goodfellow et al. (2020) Ian J. Goodfellow, Jean Pouget-Abadie, Mehdi Mirza, Bing Xu, David Warde-Farley, Sherjil Ozair, Aaron C. Courville, and Yoshua Bengio. 2020. Generative adversarial networks. Commun. ACM 63, 11 (2020), 139–144. https://doi.org/10.1145/3422622
@@
- Harper and Konstan (2016) F. Maxwell Harper and Joseph A. Konstan. 2016. The MovieLens Datasets: History and Context. ACM Trans. Interact. Intell. Syst. 5, 4 (2016), 19:1–19:19. https://doi.org/10.1145/2827872
@@
- He et al. (2020) Xiangnan He, Kuan Deng, Xiang Wang, Yan Li, Yong-Dong Zhang, and Meng Wang. 2020. LightGCN: Simplifying and Powering Graph Convolution Network for Recommendation. In Proceedings of the 43rd International ACM SIGIR conference on research and development in Information Retrieval, SIGIR 2020, Virtual Event, China, July 25-30, 2020, Jimmy X. Huang, Yi Chang, Xueqi Cheng, Jaap Kamps, Vanessa Murdock, Ji-Rong Wen, and Yiqun Liu (Eds.). ACM, 639–648. https://doi.org/10.1145/3397271.3401063
@@
- He et al. (2017) Xiangnan He, Lizi Liao, Hanwang Zhang, Liqiang Nie, Xia Hu, and Tat-Seng Chua. 2017. Neural Collaborative Filtering. In Proceedings of the 26th International Conference on World Wide Web, WWW 2017, Perth, Australia, April 3-7, 2017, Rick Barrett, Rick Cummings, Eugene Agichtein, and Evgeniy Gabrilovich (Eds.). ACM, 173–182. https://doi.org/10.1145/3038912.3052569
- Ho et al. (2020) Jonathan Ho, Ajay Jain, and Pieter Abbeel. 2020. Denoising Diffusion Probabilistic Models. In Advances in Neural Information Processing Systems 33: Annual Conference on Neural Information Processing Systems 2020, NeurIPS 2020, December 6-12, 2020, virtual, Hugo Larochelle, Marc’Aurelio Ranzato, Raia Hadsell, Maria-Florina Balcan, and Hsuan-Tien Lin (Eds.). https://proceedings.neurips.cc/paper/2020/hash/4c5bcfec8584af0d967f1ab10179ca4b-Abstract.html